# Skewed score: a statistical framework to assess autograders

### v2

@@ 1.2 Integrating the autograder assessment with the research question
Using the posterior probabilities from the fitted GLM, Florence can simulate scores and compute a distribution over $\alpha$ values, giving her not only a point estimate but also a full uncertainty interval around inter-rater agreement. This is shown in fig. 6, where the red cross denotes the standard Krippendorff $\alpha$ computed directly from the observed scores, and the blue distribution represents the GLM-based estimate. This is a significant improvement over traditional methods, which typically return a single value and do not account for sampling variability or data sparsity.
@@
However, while helpful, such agreement metrics have important limitations: they do not distinguish between disagreement due to noise and disagreement due to systematic bias. In fig. 6, both the traditional and GLM-based $\alpha$ estimates are below zero, suggesting substantial disagreement. But this does not tell us why graders disagree. To answer this question, Florence turns back to the right panel of fig. 5, where she observes that human graders have a strong positive main effect relative to autograders, meaning they consistently assign higher scores. She also notes that while items differ in average difficulty, the grader–item interaction effects are close to zero, indicating that graders do not differ much in how they score specific items.
@@
She then samples from these bias-adjusted posterior predictions and recomputes Krippendorff’s $\alpha$. The resulting agreement score posterior, shown as the green distribution in fig. 6, is substantially higher, indicating that most of the disagreement was driven by consistent shifts in scoring behaviour rather than inconsistency in judgment. Notably, the green distribution is also wider than the others. This is most likely because the model has removed a major source of predictable variation, i.e., the average bias associated with each grader, but has not explicitly accounted for individual differences in scoring behaviour. As a result, variability that was once explained by grader identity now becomes unexplained, increasing uncertainty in the prediction.
@@
Let’s imagine that Florence wants to compare the quality of outputs generated by three different LLMs. She chooses a pairwise evaluation format, where each grader (e.g., herself or an autograder) is repeatedly shown two responses to the same prompt - each generated by a different LLM - and must select the better response. An example of such data can be seen in the left panel of fig. 7. Each bar represents a pairwise comparison (e.g., “LLM A vs. LLM B”), and its height reflects how frequently the first listed model (e.g., LLM A) was chosen. To model this, we can use a binomial GLM with a logit link function. The outcome variable $y_{i}$ indicates whether the first model in the pair was chosen ($y_{i}=1$) or not ($y_{i}=0$), and we include a categorical effect in the model to denote the LLM pair being compared.
Florence’s younger brother, always up-to-date with ML controversies, recently told her that some autograders may systematically prefer longer outputs even if those outputs are not of higher quality, a phenomenon commonly referred to as length bias [Zheng et al., 2023; Dubois et al., 2024]. To capture this bias, she adds a continuous predictor capturing the token-length difference between the two outputs. As there are two graders (herself and the autograder), which might have different biases, she computes one such predictor per grader. To test the existence of the length bias formally, she compares the model with and without this term (cf. fig. 11 in appendix B). For demonstration purposes, let’s look at the model with grader-specific length bias here:
$\begin{gathered}y_{i}\sim\text{Binomial}(1,p_{i})\\ \text{logit}(p_{i})=\beta_{0}+\beta_{1}(X_{i}^{\text{LLM pair}})+\beta_{2}(X_{i}^{\text{grader}})+\beta_{3}(X_{i}^{\text{grader}})\cdot X_{i}^{\text{lengthDiff}}\\ \beta_{3}(X_{i}^{\text{grader}})\sim\mathcal{N}(\mu_{\text{lengthDiff}},\sigma_{\text{lengthDiff}}^{2})\\ \mu_{\text{lengthDiff}}\sim\mathcal{N}(0,0.5)\\ \sigma_{\text{lengthDiff}}\sim\text{HalfNormal}(1.0)\\ \end{gathered}$ (7)
@@
Inspecting the estimated effects in the right panel of fig. 7, we observe a positive effect for the grader-specific length bias parameter ($\beta_{3}$), particularly for the autograder. From this, Florence can infer that the autograder is more likely to select longer outputs, irrespective of their intrinsic quality. In other words, there appears to be an implicit association between output length and perceived correctness in the autograder’s judgments. By explicitly quantifying such biases within the model, Florence can more reliably interpret differences in LLM rankings. For example, if LLM A wins most comparisons but consistently produces longer outputs, Florence might question: “Is LLM A genuinely better, or simply more verbose? Can I really trust the autograder’s judgements?”. Additionally, by examining the estimated LLM pair parameters ($beta_{1}$), she can verify whether the observed preferences follow a consistent ranking or if there are intransitive (cyclical) patterns. Here, the estimated parameters indicate a consistent ordering: LLM A tends to be preferred to LLM B and LLM C, and LLM B tends to be preferred to LLM C. This integrated statistical framework empowers her to disentangle and quantify these systematic biases and assess preference consistency, leading to deeper and more reliable conclusions.
@@ Conclusion
The examples presented are by no means exhaustive. Many other applications and extensions of GLMs are possible, and we hope this work provides a clear and practical starting point for researchers seeking to adapt the framework to their own evaluation scenarios. To support practical adoption, we have summarised common evaluation questions and their implementation in a GLM setup in table 1. All models presented in this paper are implemented in the open-source HiBayes package, and reproducible notebooks for all examples are available in a public repository.
@@ Appendix B Supplementary figures: Model comparisons
Figure 8: Model comparison for the statistical analysis in Question 1.1 (quantifying the mean difference between scores assigned by an autograder and a human grader). Left panel: Leave-One-Out cross-validation (LOO) scores. Right panel: Widely Applicable Information Criterion (WAIC). Both metrics approximate the Expected Log Predictive Density (ELPD), a measure of predictive accuracy (higher values indicate better performance). Comparing these models helps determine whether including a grader-effect term (eq. 2) is justified by improved predictive performance. In this case, the model with the grader effect clearly outperforms the null model, supporting closer examination of the grader’s impact on scoring
Figure 9: Model comparison for the statistical analysis in Question 2 (Do autograders favour their own generation?). Left panel: Leave-One-Out cross-validation (LOO) scores. Right panel: Widely Applicable Information Criterion (WAIC). Both metrics approximate the Expected Log Predictive Density (ELPD), a measure of predictive accuracy (higher values indicate better performance). Comparing models helps determine whether the added complexity of including an interaction term is justified by improved predictive performance. In this case, the model with the grader–LLM interaction (eq. 4) performs slightly better than the model without interaction, supporting a closer examination of potential self-bias effects.
Figure 10: Model comparison for the statistical analysis in Question 3 (Do autograders differ systematically from human experts?). Left panel: Leave-One-Out cross-validation (LOO) scores. Right panel: Widely Applicable Information Criterion (WAIC). Both metrics approximate the Expected Log Predictive Density (ELPD), a measure of predictive accuracy (higher values indicate better performance). The models perform similarly, which is expected given that the data is simulated without an explicitly hierarchical structure. Here we choose the hierarchical model (eq. 5) to demonstrate how to interpret its parameters. In practice, when models perform similarly, researchers should favour the simpler model unless theoretical or interpretability considerations justify the added complexity.
Figure 11: Model comparison for the statistical analysis in Question 5 (Do autograders favour longer outputs?). Left panel: Leave-One-Out cross-validation (LOO) scores. Right panel: Widely Applicable Information Criterion (WAIC). Both metrics approximate the Expected Log Predictive Density (ELPD), a measure of predictive accuracy (higher values indicate better performance). Comparing these models tests whether including a length-effect term (eq. 7) significantly improves predictive performance. The model including the length effect clearly outperforms the simpler model without it, justifying a closer investigation into grader-specific length biases.
@@ Appendix C Supplementary question: Is the grading scale well calibrated?
In this section we explore the grading process, and why we chose an ordered logistic regression model. To ensure consistent evaluation, Florence needs to establish standardised scoring across all graders. One simple approach would be to instruct everyone (humans and autograders) to count the number of relevant keywords from a predefined list. Besides obvious issues (e.g., synonyms need to be accounted for), this approach is not scalable as it require building explicit keyword lists for each open-ended question. Florence, as often done in practice, will instead develop a grading rubric that can be applied to each question. She might come up with something like in table 2.
@@
Going back to Florence, to better capture differences in model capabilities, she decides to examine her grading scale. To do this, she can look at the learned cutpoint parameters of the ordered logistic regression. Taking eq. 3 for example, we can write the cumulative probability more explicitly as:
@@
Once Florence fits this model, she can analyse the inferred cutpoint values. Those can be found in table 3. She observes that the distance c1-c2, and c2-c3, is below 1 unit, whereas the distance c6-c7 and c7-c8 is above 1.5 units. This large jump suggests that her scale lacks sensitivity around c6-c8 (i.e., many performance scores are clustering there). From this, she could decide that she wants to better capture difference in that region and therefore adds some intermediate categories in this area.

### v3

@@ 1.2 Integrating the autograder assessment with the research question
Using the posterior probabilities from the fitted GLM, Florence can simulate scores and compute a distribution over $\alpha$ values, giving her not only a point estimate but also a full uncertainty interval around inter-rater agreement. This is shown in fig.˜6, where the red cross denotes the standard Krippendorff $\alpha$ computed directly from the observed scores, and the blue distribution represents the GLM-based estimate. This is a significant improvement over traditional methods, which typically return a single value and do not account for sampling variability or data sparsity.
@@
However, while helpful, such agreement metrics have important limitations: they do not distinguish between disagreement due to noise and disagreement due to systematic bias. In fig.˜6, both the traditional and GLM-based $\alpha$ estimates are below zero, suggesting substantial disagreement. But this does not tell us why graders disagree. To answer this question, Florence turns back to the right panel of fig.˜5, where she observes that human graders have a strong positive main effect relative to autograders, meaning they consistently assign higher scores. She also notes that while items differ in average difficulty, the grader–item interaction effects are close to zero, indicating that graders do not differ much in how they score specific items.
@@
She then samples from these bias-adjusted posterior predictions and recomputes Krippendorff’s $\alpha$. The resulting agreement score posterior, shown as the green distribution in fig.˜6, is substantially higher, indicating that most of the disagreement was driven by consistent shifts in scoring behaviour rather than inconsistency in judgment. Notably, the green distribution is also wider than the others. This is most likely because the model has removed a major source of predictable variation, i.e., the average bias associated with each grader, but has not explicitly accounted for individual differences in scoring behaviour. As a result, variability that was once explained by grader identity now becomes unexplained, increasing uncertainty in the prediction.
@@
Let’s imagine that Florence wants to compare the quality of outputs generated by three different LLMs. She chooses a pairwise evaluation format, where each grader (e.g., herself or an autograder) is repeatedly shown two responses to the same prompt - each generated by a different LLM - and must select the better response. An example of such data can be seen in the left panel of fig.˜7. Each bar represents a pairwise comparison (e.g., “LLM A vs. LLM B”), and its height reflects how frequently the first listed model (e.g., LLM A) was chosen. To model this, we can use a binomial GLM with a logit link function. The outcome variable $y_{i}$ indicates whether the first model in the pair was chosen ($y_{i}=1$) or not ($y_{i}=0$), and we include a categorical effect in the model to denote the LLM pair being compared.
Florence’s younger brother, always up-to-date with ML controversies, recently told her that some autograders may systematically prefer longer outputs even if those outputs are not of higher quality, a phenomenon commonly referred to as length bias [Zheng et al., 2023; Dubois et al., 2024]. To capture this bias, she adds a continuous predictor capturing the token-length difference between the two outputs. As there are two graders (herself and the autograder), which might have different biases, she computes one such predictor per grader. To test the existence of the length bias formally, she compares the model with and without this term (cf. fig.˜11 in appendix˜B). For demonstration purposes, let’s look at the model with grader-specific length bias here:
[…]
@@
Inspecting the estimated effects in the right panel of fig.˜7, we observe a positive effect for the grader-specific length bias parameter ($\beta_{3}$), particularly for the autograder. From this, Florence can infer that the autograder is more likely to select longer outputs, irrespective of their intrinsic quality. In other words, there appears to be an implicit association between output length and perceived correctness in the autograder’s judgments. By explicitly quantifying such biases within the model, Florence can more reliably interpret differences in LLM rankings. For example, if LLM A wins most comparisons but consistently produces longer outputs, Florence might question: “Is LLM A genuinely better, or simply more verbose? Can I really trust the autograder’s judgements?”. Additionally, by examining the estimated LLM pair parameters ($beta_{1}$), she can verify whether the observed preferences follow a consistent ranking or if there are intransitive (cyclical) patterns. Here, the estimated parameters indicate a consistent ordering: LLM A tends to be preferred to LLM B and LLM C, and LLM B tends to be preferred to LLM C. This integrated statistical framework empowers her to disentangle and quantify these systematic biases and assess preference consistency, leading to deeper and more reliable conclusions.
@@ Conclusion
The examples presented are by no means exhaustive. Many other applications and extensions of GLMs are possible, and we hope this work provides a clear and practical starting point for researchers seeking to adapt the framework to their own evaluation scenarios. To support practical adoption, we have summarised common evaluation questions and their implementation in a GLM setup in table˜1. All models presented in this paper are implemented in the open-source HiBayes package, and reproducible notebooks for all examples are available in a public repository.
@@ Appendix B Supplementary figures: Model comparisons
Figure 8: Model comparison for the statistical analysis in Question 1.1 (quantifying the mean difference between scores assigned by an autograder and a human grader). Left panel: Leave-One-Out cross-validation (LOO) scores. Right panel: Widely Applicable Information Criterion (WAIC). Both metrics approximate the Expected Log Predictive Density (ELPD), a measure of predictive accuracy (higher values indicate better performance). Comparing these models helps determine whether including a grader-effect term (eq.˜2) is justified by improved predictive performance. In this case, the model with the grader effect clearly outperforms the null model, supporting closer examination of the grader’s impact on scoring
Figure 9: Model comparison for the statistical analysis in Question 2 (Do autograders favour their own generation?). Left panel: Leave-One-Out cross-validation (LOO) scores. Right panel: Widely Applicable Information Criterion (WAIC). Both metrics approximate the Expected Log Predictive Density (ELPD), a measure of predictive accuracy (higher values indicate better performance). Comparing models helps determine whether the added complexity of including an interaction term is justified by improved predictive performance. In this case, the model with the grader–LLM interaction (eq.˜4) performs slightly better than the model without interaction, supporting a closer examination of potential self-bias effects.
Figure 10: Model comparison for the statistical analysis in Question 3 (Do autograders differ systematically from human experts?). Left panel: Leave-One-Out cross-validation (LOO) scores. Right panel: Widely Applicable Information Criterion (WAIC). Both metrics approximate the Expected Log Predictive Density (ELPD), a measure of predictive accuracy (higher values indicate better performance). The models perform similarly, which is expected given that the data is simulated without an explicitly hierarchical structure. Here we choose the hierarchical model (eq.˜5) to demonstrate how to interpret its parameters. In practice, when models perform similarly, researchers should favour the simpler model unless theoretical or interpretability considerations justify the added complexity.
Figure 11: Model comparison for the statistical analysis in Question 5 (Do autograders favour longer outputs?). Left panel: Leave-One-Out cross-validation (LOO) scores. Right panel: Widely Applicable Information Criterion (WAIC). Both metrics approximate the Expected Log Predictive Density (ELPD), a measure of predictive accuracy (higher values indicate better performance). Comparing these models tests whether including a length-effect term (eq.˜7) significantly improves predictive performance. The model including the length effect clearly outperforms the simpler model without it, justifying a closer investigation into grader-specific length biases.
@@ Appendix C Supplementary question: Is the grading scale well calibrated?
In this section we explore the grading process, and why we chose an ordered logistic regression model. To ensure consistent evaluation, Florence needs to establish standardised scoring across all graders. One simple approach would be to instruct everyone (humans and autograders) to count the number of relevant keywords from a predefined list. Besides obvious issues (e.g., synonyms need to be accounted for), this approach is not scalable as it require building explicit keyword lists for each open-ended question. Florence, as often done in practice, will instead develop a grading rubric that can be applied to each question. She might come up with something like in table˜2.
@@
Going back to Florence, to better capture differences in model capabilities, she decides to examine her grading scale. To do this, she can look at the learned cutpoint parameters of the ordered logistic regression. Taking eq.˜3 for example, we can write the cumulative probability more explicitly as:
@@
Once Florence fits this model, she can analyse the inferred cutpoint values. Those can be found in table˜3. She observes that the distance c1-c2, and c2-c3, is below 1 unit, whereas the distance c6-c7 and c7-c8 is above 1.5 units. This large jump suggests that her scale lacks sensitivity around c6-c8 (i.e., many performance scores are clustering there). From this, she could decide that she wants to better capture difference in that region and therefore adds some intermediate categories in this area.